OpenStudy (anonymous):

Solve for z in z^2-i = 0?

OpenStudy (anonymous):

I got to z = sqrt(i), at which point I substituted for sqrt(i)=a + bi. Solving gave me \[\frac{ \sqrt(2) }{ 4 } + \frac{ 1 }{ \sqrt{2} }i\]

OpenStudy (anonymous):

But the book says the answer is

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{2} }-\frac{ 1 }{ \sqrt{2} }i\]

OpenStudy (anonymous):

No wait, make that a plus

TheSmartOne (thesmartone):

\[i=\sqrt{-1}\]

TheSmartOne (thesmartone):

that is all I can help you with.. I haven't taken Calculus yet...

OpenStudy (anonymous):

don't answer questions you can't answer...

TheSmartOne (thesmartone):

sorry.. I didn't see it was labeled under calculus.. Any question posted in the sub categories of math jsut show up in math...

OpenStudy (anonymous):

\[\begin{align*} z^2-i&=0\\\\ z^2&=i\\\\ z&=i^{1/2}\\\\ z&=\left(\cos\frac{\pi}{2}+i\sin\frac{\pi}{2}\right)^{1/2}\\\\ z&=\begin{cases} \cos\dfrac{\dfrac{\pi}{2}}{2}+i\sin\dfrac{\dfrac{\pi}{2}}{2}\\\\ \cos\left(\dfrac{\dfrac{\pi}{2}+\pi}{2}\right)+i\sin\left(\dfrac{\dfrac{\pi}{2}+\pi}{2}\right) \end{cases} \end{align*}\]

OpenStudy (anonymous):

Hold on, that second root doesn't seem to be right... It's giving me \(-i\) when I square it.

OpenStudy (anonymous):

@SithsAndGiggles the second should be pi/4 + pi

OpenStudy (anonymous):

5pi/4

OpenStudy (anonymous):

Ah yes, the numerator in the argument should be \(\dfrac{\pi}{2}+2\pi\)... Thanks

OpenStudy (anonymous):

Right, I forgot I could convert to polar form... thanks

OpenStudy (anonymous):

yw